Find the effective annual interest rate of an account containing $3000 which is compounded continuously at 8%
Answer by stanbon(75887)

Find the effective annual interest rate of an account containing $3000 which is compounded continuously at 8%
-----
Formula: A(t) = P(e^(rt))
----
e^(rt) = e^0.08 = 1.0833
==============================
Effective annual interest rate: 8.33%
=========================================
Cheers,
Stan H.
=========